120PERCENT OF A IS EQUAL TO 80 PERCENT OF B, FIND A+B

Answer by stanbon(75887)

120 PERCENT OF A IS EQUAL TO 80 PERCENT OF B, FIND A+B
-----
Equation:
1.2A = 0.8B
A = [0.8/1.2]B
A = (2/3)B
-------
A + B = (2/3)B + B = (5/3)B
==========================
cheers,
Stan H.
=================